Scott is four times older than Beth. In six years from now Scott will only be three times older than Beth. Write down equation to describe the relationship between their ages. Now solve the equation to find out how old each of them is now
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Scott is four times older than Beth. In six years from now Scott will only be three times older than Beth. Write down equation to describe the relationship between their ages. Now solve the equation to find out how old each of them is now
----
Now:
S = 5B
In 6 years:
S + 6 = 3*(B + 6)
===================
Sub for C
5B + 6 = 3B + 18
2B = 12
B = 6
S = 30

Answer by ikleyn(52817)

.

            In English, "four times older than . . . " means "five times as old as . . . ".

            Also, "three times older than . . . " means "four times as old as . . . ".

            Therefore, the correct solution is as follows:


 
S = 5B.                    (1)
S + 6 = 4*(B+6).           (2)


It is your starting system of two equations, which is the literal translation of the condition to Math.


To solve the system, substitute (1) into equation (2), replacing S. You will get a single equation for B


5B + 6 = 4*(B+6).


Simplify and solve it for B:


5B + 6 = 4B + 24

B + 6 = 24

B = 24 - 6 = 18.


Answer.  Beth is 18 years old.   Scott is 5*B = 5*18 = 90 years old.

Solved.

------------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.